OpenStudy (anonymous):

Using a directrix of y = 5 and a focus of (4, 1), what quadratic function is created?

OpenStudy (anonymous):

f(x) = one fourth (x - 4)2 - 3 f(x) = one eighth (x + 4)2 - 3 f(x) = -one eighth (x - 4)2 + 3 f(x) = -one fourth (x + 4)2 - 3

OpenStudy (ranga):

The definition of a parabola is that any point on the parabola is the same distance from the focus and the directrix. Let (x,y) be a point on the parabola. Its distance from the focus (or distance squared) = (x - 4)^2 + (y -1)^2 The squared of the distance between the point (x,y) and y = 5 is (y - 5)^2 Equate the two distances (squared): (x - 4)^2 + (y -1)^2 = (y - 5)^2 Take the y term to RHS: (x - 4)^2 = (y - 5)^2 - (y -1)^2 = (y - 5 + y - 1)(y - 5 - y + 1) (since a^2 - b^2 = (a+b)(a-b)) = (2y-6)(-4) = 2(y-3)(-4) =-8(y-3) -1/8(x-4)^2 = y - 3 y = -1/8(x-4)^2 + 3 f(x) = -1/8(x-4)^2 + 3

OpenStudy (anonymous):

Thank you again ranga, and if you could also help me with another? Solve 2x^2 + 16x + 34 = 0 i got x= -4+i and x=-4-i is that the same as (-4 +or- i) or is it the same as (x-4 +or-i) or is it neither?

OpenStudy (ranga):

2x^2 + 16x + 34 = 0 factor out 2 2(x^2 + 8x + 17) = 0 x^2 + 8x + 17 = 0 x = [ -8 +/- sqrt { (-8)^2 - (4)(1)(17) } ] / 2 = { -8 +/- sqrt(-4) } / 2 = -4 +/- i Yes. -4 +/- i is shorthand representation of two solutions: -4 + i and -4 - i

OpenStudy (anonymous):

thank you so much for your help and explanations!!!

OpenStudy (ranga):

You are very welcome.